How can Planck's constant be deduced from the graph of stopping potential vs frequency of light ? How can the work function be obtained from this graph?
Answer #1

slope = h/e

==> h = slope*e (here e = 1.6*10^-19 C)

y- intercept = -wo/e

==> wo = y-intercept*e  (here e = 1.6*10^-19 C)

Stopping Potential Vs Slope h/e 0 Frequency f